OpenStudy (anonymous):

dhattraditya help me please! A circular must deliver 18 watts of energy .If each square centimter of the cell that is in sunlight produces 0.01 watt of energy, how long must the radius of the cells? Page 506# 76

OpenStudy (anonymous):

Can you just say that you need 18/.01 square centimeters to produce 18 watts? so 1,800 square centimeters. Maybe you can use he formula for Area to give you the answer for radius? \[\pi(r^2)=1,800\] so \[r^2=1,800/\pi\] then r = 23.93653682

OpenStudy (anonymous):

just a thought

OpenStudy (anonymous):

ty